OpenStudy (anonymous):

find the difference quotient for 6/(x^2)

OpenStudy (amistre64):

???

jimthompson5910 (jim_thompson5910):

he wants to find \[\large \frac{f(x+h)-f(x)}{h}\] where in this case \[\large f(x)=\frac{6}{x^2}\]

OpenStudy (anonymous):

yeah i dont know what to do once i have \[(6/(x^2+2xh+h^2) -6/x^2)/h\]

jimthompson5910 (jim_thompson5910):

multiply EVERY term by the inner LCD \[\large x^2(x^2+2xh+h^2)\] this will clear out the inner fractions.

OpenStudy (anonymous):

so then you get (6x^2(x^2+2xh+h^2)-6x^2(x^2+2xh+h^2))/h?

jimthompson5910 (jim_thompson5910):

not quite, it will give you \[\large \frac{6x^2-6(x^2+2xh+h^2)}{hx^2(x^2+2xh+h^2)}\]

jimthompson5910 (jim_thompson5910):

what's next then?

OpenStudy (anonymous):

0/h?

jimthompson5910 (jim_thompson5910):

distribute the 6 through to get \[\large \frac{6x^2-6x^2-12xh-6h^2}{hx^2(x^2+2xh+h^2)}\] notice the 6x^2 terms cancel, so we then get \[\large \frac{-12xh-6h^2}{hx^2(x^2+2xh+h^2)}\] now factor out an h from the numerator \[\large \frac{h(-12x-6h)}{hx^2(x^2+2xh+h^2)}\] this h up top cancels with the h down below to give us \[\large \frac{-12x-6h}{x^2(x^2+2xh+h^2)}\] From there, we can distribute the terms in the bottom to give us \[\large \frac{-12x-6h}{x^4+2x^3h+x^2h^2}\] and this is as far as we go since we can't do anything with h (not yet at least)

OpenStudy (anonymous):

i dont know how to move on

jimthompson5910 (jim_thompson5910):

that's the last step

OpenStudy (anonymous):

i dont know how to simplify though cant it be simplified?

jimthompson5910 (jim_thompson5910):

if you're familiar with limits, you would then plug in h=0 and simplify to get \[\large -\frac{12}{x^3}\]

jimthompson5910 (jim_thompson5910):

but since I'm assuming you're not, the last step is then \[\large \frac{-12x-6h}{x^4+2x^3h+x^2h^2}\]

OpenStudy (anonymous):

alright thanks